Please help me solve this question. I dont seem to get it. Help is appreciated.
If the graph of the quadratic function y=ax+3 must pass through the point (4100)then what is the value of a to the nearest hundredths
I have no working to show because I havent got a clue about this.
Thanks 4 helping out.

Answer by stanbon(75887)

If the graph of the quadratic function y=ax^2+3 must pass through the point (4,100)then what is the value of a to the nearest hundredthstu
---
Substitute for x and y and solve for "a":
100 = a(4^2)+3
97 = a*16
a = 6.06
=============
Cheers,
Stan H.
=============